# TDAvec: Computing Vector Summaries of Persistence Diagrams for Topological Data Analysis in R and Python

### v2

@@ I Introduction
The theory of persistent homology is one of the popular tools in topological data analysis (TDA) to analyze data with underlying shape structure [Car09, EH10, CM21]. In this context, a single data observation could be a collection of points lying in a metric space, an image, a graph or a time series. The basic idea behind persistent homology is to build a nested sequence (or filtration) of simplicial complexes (indexed by a scale parameter) on top of data points and keep a record of the appearance and disappearance of various topological features at different scale values. Here, these topological features are ”holes” of different dimensions – connected components, loops, voids, and their higher-dimensional versions whose emergence and subsequent disappearance are tracked using a concept of homology from algebraic topology. From a geometric point of view, simplicial complexes consist of vertices, edges, triangles, tetrahedra etc., glued together and serve as a means for recovering (at least partially) the underlying shape information which is lost during sampling [NS13].
A topological descriptor outputted by the persistent homology encoding the shape of data is called a persistence diagram (PD). Mathematically, a $k$-dimensional PD is a multi-set of points $D=\{(b_{i},d_{i})\}_{i=1}^{N}$, where each point $(b_{i},d_{i})$ corresponds to a topological feature of homological dimension $k$ (0 if a connected component, 1 if a loop, 2 if a void, etc) with the $x$-coordinate representing the scale at which this feature is born (or created), and the $y$-coordinate representing the scale at which it dies (or disappears). In practice, one is usually interested in applying a machine learning method to PDs to make further inferences from data. However, the fact that PDs do not form a Hilbert space, which is a feature (or an input) space for a wide class of machine learning methods, limits their direct of use in applications. To overcome this challenge, kernel methods and vectorization techniques are commonly used [CL22]. The kernel approach involves defining a notion of similarity between pairs of PDs, whereas the vectorization methods aim to transform PDs into finite-dimensional feature vectors that can be used as input for many standard machine learning models. Such vector summaries of PDs are computed in two steps: first one constructs a respective summary function from a given PD and then vectorizes it using either one or two dimensional grid of scale values. In recent years, the kernel and vectorization approaches have proven successful and gained prominence in the applied TDA literature (see [HMR21] for a survey of applications of TDA in machine learning).
@@
The rest of the paper is organized as follows. In the next section we describe in more details our package and prove its necessity. In the next section a short overview of existing R and python packages is presented. In section III a short description of our package is given, followed by some usage examples. Last section is n is reserved for discussion. In the appendix you will find definitions of all used in our work vectorization summaries.
@@ II-A R tools for TDA
The computational tools for TDA in the R environment are provided through various packages [^1] In this section, we only focus on R packages for TDA that are available on the CRAN repository. such as TDA [FKL+21], TDAstats [WWDS18], kernelTDA [PP20], TDAmapper [PMS15], TDAkit [YY21], tdaunif [BDC24], TDApplied [BF24] and ripserr [WPS20] (see Table I for an overview of these packages in terms of their scope and areas of focus).
@@
The TDA package is the largest R package for TDA. The TDA package offers tools to compute PDs for commonly used types of filtrations such as Vietoris-Rips, Alpha and Alpha shape. It also allows to construct more general sublevel set filtrations and compute the corresponding PDs. Moreover, the TDA package provides implementations to plot PDs and compute bottleneck and Wasserstein distances between them. TDAstats offers a variety of tools for conducting statistical inference (such as hypothesis testing) on PDs. Compared to the TDA package, it computes PDs much faster for Vietoris-Rips filtrations based on the Ripser C++ library [@Bauer2021Ripser] and offers more aesthetic visualization of the diagrams using the ggplot2 package [@ggplot2]. The kernelTDA package contains implementations of popular kernel-based methods for TDA such as geodesic Gaussian kernel, geodesic Laplacian kernel, persistence Fisher kernel and persistence sliced Wasserstein kernel. For computing the Wasserstein distance between a pair of PDs, unlike the TDA package, it uses an iterative procedure to reasonably approximate the exact distance which that leads to a considerable reduction in run-time cost. The ripserr package allows a fast computation of PDs for filtrations on Vietoris-Rips and cubical complexes using the Ripser C++ library. The TDApplied and TDAkit packages provides various tools to integrate topological features (PDs or their vector summaries) into machine and statistical learning settings. The tdaunif is a useful package if one needs to sample points from various manifolds such as a klein-bottle, an ellipse or a torus. TDAmapper offers tools to visualize high-dimensional data by constructing the so-called Mapper graphs that preserve its topological structure.
@@ II-B Python tools for TDA
Several Python libraries are available for TDA, such as Giotto-tda [TLT+21], Ripser [CNR18], Gudhi [Rou20], Dionysus 2 [Mor18], Scikit-tda [ST19] which is a container library for such projects as Ripser, Persim [Sau19], KeplerMapper [VVSEM19], etc. Giotto-tda is a powerful library that integrates with the popular machine learning library scikit-learn, offering tools for persistent homology and visualizations of persistence diagrams. Ripser focuses on fast computation of Vietoris-Rips complexes, especially for large datasets. Gudhi provides a wide range of topological tools for simplicial complexes, persistent homology, and topological signatures. Dionysus 2 offers fast computation of persistent homology and cohomology, with an emphasis on flexibility and efficiency. Persim focuses on tools for working with PDs. It contains implementations of commonly used vectorization and kernel methods for PDs. KeplerMapper implements the TDA Mapper algorithm to visualize high-dimensional data. Scikit-tda is another package that integrates with scikit-learn, simplifying the application of TDA to typical machine learning tasks.
For a more comprehensive list of Python libraries for TDA and their functionality, we refer the readers to [awe24].
@@ III-A Statement of need and Short Package Description
1. It expands the list of implemented vector summaries of PDs by providing vectorization of eight functional summaries found in the TDA literature: Betti function, persistence landscape function, persistence silhouette function, persistent entropy summary function [AGDST20], Euler characteristic curve [RW14], normalized life curve [CL22], persistence surface [AEK+17] and persistence block [CILS22]. In addition, in the updated version o the article such vectorizations as simple simple statistics, Tropical Coordinates, and Template Functions [PMK23] were added. In the Appendeix you an see definitions of theese vetorizations.
2. A univariate summary function $f$ of a PD is typically vectorized by evaluating it at each point of a superimposed one dimensional grid and arranging the resulting values into a vector:
@@
[…]
@@
3. To achieve higher computational efficiency, all code behind the vector summaries of TDAvec is written in C++. For example, in R, computing the persistence landscape from a PD with the TDAvec package is more than 200 times faster than with the TDA package.
The TDAvec R package and a vignette showing its basic usage with examples are available on the CRAN repository [^2] https://cran.r-project.org/web/packages/TDAvec/index.html. For Python examples, we refer the readers to sample notebook presented in [LI24].
@@ III-B Usage Examples
Using the created data frame, we can easily compare performance accuracies (see table III). It’s also simple to check the accuracy of the models by inspecting true vs. predicted scatter plots (see figure).
@@
TABLE III: Test Accuracied of different ML methods
@@ Appendix A Definitions of the summary functions in TDAvec
where $\beta_{0},\beta_{1},\ldots,\beta_{d}$ are the Betti curves corresponding to persistence diagrams $D_{0},D_{1},\ldots,D_{d}$ of dimensions $0,1,\ldots,d$ respectively, all computed from the same filtration. Note that unlike all other vectorizations the Euler characteristics curve is calculated using all available dimenstions.
@@
9) Functional Data Analysis:
$g(t)=\sum_{i}I_{[b_{i},d_{i})]}(t)=\sum_{k}c_{k}\xi_{k}(t)$
10) Simple statistics:
For a given persistence diagram $D=\{(b_{i},d_{i})\}_{i=1}^{N}$ (corresponding to a specified homological dimension), computeStats() calculates descriptive statistics of the birth, death, midpoint (the average of birth and death), and lifespan (death minus birth) values. Additionally, it computes the total number of points and entropy of the lifespan values. Points in $D$ with infinite death values are ignored.
11) Tropical Coordinates:
This vectorization for each homologial dimension computes the following seven tropical coordinates based on the lifespans (or persistence) $\lambda_{i}=d_{i}-b_{i}$:
@@
    12) Template Function Vectorization
    For a given persistence diagram $D=\{(b_{i},d_{i})\}_{i=1}^{N}$ (corresponding to a specified homological dimension), this vectorization uses a collection of tent template functions defined by

### v3

# Vectorization of Persistence Diagrams for Topological Data Analysis in R and Python Using TDAvec Package

@@ I Introduction
Persistent homology has emerged as one of the most powerful tools in topological data analysis (TDA) for uncovering the geometric and topological structure of complex data [7, 12, 10]. Within this framework, a data observation can take many forms—such as a point cloud in a metric space, an image, a graph, or a time series. The core idea of persistent homology is to construct a nested sequence (or filtration) of simplicial complexes indexed by a scale parameter, and to track how topological features appear and disappear across scales. These features correspond to “holes” of different dimensions—connected components, loops, voids, and their higher-dimensional analogues—captured through the algebraic machinery of homology. From a geometric perspective, simplicial complexes, composed of vertices, edges, triangles, and higher-dimensional simplices, provide a discrete representation that partially recovers the shape of the underlying space from sampled data [17].
The principal output of persistent homology is the persistence diagram (PD), a concise topological summary that encodes the lifetime of features across scales. A $k$-dimensional PD is defined as a multiset of points $D={(b_{i},d_{i})}_{i=1}^{N}$, where each point $(b_{i},d_{i})$ represents a feature of homological dimension kk—with $k=0$ for connected components, $k=1$ for loops, and $k=2$ for voids. The birth coordinate $b_{i}$ indicates the scale at which a feature appears, and the death coordinate $d_{i}$ the scale at which it disappears.
In many practical applications, persistence diagrams serve as inputs to machine learning and statistical models. However, since PDs do not form a Hilbert space, they cannot be directly used with most standard algorithms that rely on vector or inner-product representations. To overcome this limitation, researchers have developed two major strategies: kernel methods and vectorization techniques [9]. Kernel methods define similarity measures between pairs of PDs, while vectorization approaches transform diagrams into finite-dimensional feature vectors that can be readily processed by machine learning models. In the latter case, the procedure typically involves constructing a summary function from a PD—such as a persistence image, landscape, or entropy profile—and then discretizing it over a one- or two-dimensional grid of scale values. Both paradigms have been widely adopted and proven effective in applied TDA and data-driven research [14].
Despite the growing interest in topological data analysis, software support for vectorization of persistence diagrams remains fragmented across languages and packages, often differing in implementation details, computational efficiency, and interface design. To address this gap, we introduce TDAvec, a unified and efficient library for computing vector summaries of persistence diagrams in both R and Python. The package provides a consistent interface for several widely used vectorization techniques, allowing researchers to seamlessly integrate topological descriptors into machine learning workflows
@@
The rest of the paper is organized as follows. In the next section we describe in more details our package and prove its necessity. In the next section a short overview of existing R and python packages is presented. In section III a short description of our package is given, followed by some usage examples. Last section is reserved for discussion. In the appendix you will find definitions of all used in our work vectorization summaries.
@@ II-A R tools for TDA
The computational tools for TDA in the R environment are provided through various packages [^1] In this section, we only focus on R packages for TDA that are available on the CRAN repository. such as TDA [13], TDAstats [29], kernelTDA [20], TDAmapper [19], TDAkit [30], tdaunif [5], TDApplied [6] and ripserr [28] (see Table I for an overview of these packages in terms of their scope and areas of focus).
@@
The TDA package is the largest R package for TDA. The TDA package offers tools to compute PDs for commonly used types of filtrations such as Vietoris-Rips, Alpha and Alpha shape. It also allows to construct more general sublevel set filtrations and compute the corresponding PDs. Moreover, the TDA package provides implementations to plot PDs and compute bottleneck and Wasserstein distances between them. TDAstats offers a variety of tools for conducting statistical inference (such as hypothesis testing) on PDs. Compared to the TDA package, it computes PDs much faster for Vietoris-Rips filtrations based on the Ripser C++ library [4] and offers more aesthetic visualization of the diagrams using the ggplot2 package [27]. The kernelTDA package contains implementations of popular kernel-based methods for TDA such as geodesic Gaussian kernel, geodesic Laplacian kernel, persistence Fisher kernel and persistence sliced Wasserstein kernel. For computing the Wasserstein distance between a pair of PDs, unlike the TDA package, it uses an iterative procedure to reasonably approximate the exact distance which that leads to a considerable reduction in run-time cost. The ripserr package allows a fast computation of PDs for filtrations on Vietoris-Rips and cubical complexes using the Ripser C++ library. The TDApplied and TDAkit packages provides various tools to integrate topological features (PDs or their vector summaries) into machine and statistical learning settings. The tdaunif is a useful package if one needs to sample points from various manifolds such as a klein-bottle, an ellipse or a torus. TDAmapper offers tools to visualize high-dimensional data by constructing the so-called Mapper graphs that preserve its topological structure.
@@ II-B Python tools for TDA
Several Python libraries are available for TDA, such as Giotto-tda [25], Ripser [11], Gudhi [21], Dionysus 2 [16], Scikit-tda [24] which is a container library for such projects as Ripser, Persim [23], KeplerMapper [26], etc. Giotto-tda is a powerful library that integrates with the popular machine learning library scikit-learn, offering tools for persistent homology and visualizations of persistence diagrams. Ripser focuses on fast computation of Vietoris-Rips complexes, especially for large datasets. Gudhi provides a wide range of topological tools for simplicial complexes, persistent homology, and topological signatures. Dionysus 2 offers fast computation of persistent homology and cohomology, with an emphasis on flexibility and efficiency. Persim focuses on tools for working with PDs. It contains implementations of commonly used vectorization and kernel methods for PDs. KeplerMapper implements the TDA Mapper algorithm to visualize high-dimensional data. Scikit-tda is another package that integrates with scikit-learn, simplifying the application of TDA to typical machine learning tasks.
For a more comprehensive list of Python libraries for TDA and their functionality, we refer the readers to [3].
@@ III-A Statement of need and Short Package Description
  1. 1.
    It expands the list of implemented vector summaries of PDs by providing vectorization of eight functional summaries found in the TDA literature: Betti function, persistence landscape function, persistence silhouette function, persistent entropy summary function [2], Euler characteristic curve [22], normalized life curve [9], persistence surface [1] and persistence block [8]. In addition, in the updated version o the article such vectorizations as simple statistics, Tropical Coordinates, and Template Functions [18] were added. In the Appendix you an see definitions of these vectorizations.
  2. 2.
    A univariate summary function $f$ of a PD is typically vectorized by evaluating it at each point of a superimposed one dimensional grid and arranging the resulting values into a vector:
@@
    $\Big(\frac{1}{\Delta t_{1}}\int_{t_{1}}^{t_{2}}f(t)dt,\frac{1}{\Delta t_{2}}\int_{t_{2}}^{t_{3}}f(t)dt,\ldots,\frac{1}{\Delta t_{n-1}}\int_{t_{n-1}}^{t_{n}}f(t)dt\Big)\in{R}^{n-1},$ (2)
@@
  3. 3.
    To achieve higher computational efficiency, all code behind the vector summaries of TDAvec is written in C++. For example, in R, computing the persistence landscape from a PD with the TDAvec package is more than 200 times faster than with the TDA package.
  4. 4.
    We created a Python library implementing all TDAvec functions, thus extending their availability to the Python ecosystem. The library offers the exact same interface, and a scikit-learn-type interface is also provided for completeness.
The TDAvec R package and a vignette showing its basic usage with examples are available on the CRAN repository [^2] https://cran.r-project.org/web/packages/TDAvec/index.html. For Python examples, we refer the readers to sample notebook presented in [15].
@@ III-B Usage Examples
Using the created data frame, we can easily compare performance accuracies (see table III). It’s also simple to check the accuracy of the models by inspecting true vs. predicted scatter plots (see figure 3).
Figure 3: Linear Regression comparison
@@
TABLE III: Test Accuracies of different ML methods
@@ Appendix A Definitions of the summary functions in TDAvec
where $\beta_{0},\beta_{1},\ldots,\beta_{d}$ are the Betti curves corresponding to persistence diagrams $D_{0},D_{1},\ldots,D_{d}$ of dimensions $0,1,\ldots,d$ respectively, all computed from the same filtration. Note that unlike all other vectorizations the Euler characteristics curve is calculated using all available dimensions.
@@
9) Simple statistics:
For a given persistence diagram computeStats() calculates descriptive statistics of the birth, death, midpoint (the average of birth and death), and lifespan (death minus birth) values corresponding to a specified homological dimension. Additionally, it computes the total number of points and entropy of the lifespan values. Points in $D$ with infinite death values are ignored.
10) Tropical Coordinates:
This vectorization for each homological dimension computes the following seven tropical coordinates based on the lifespans (or persistence) $\lambda_{i}=d_{i}-b_{i}$:
@@
    11) Template Function Vectorization
    This vectorization uses a collection of tent template functions defined by
@@
    12) Algebraic Functions
    This vectorization computes following four algebraic functions bases on birth and death values:
    $f_{1}$ $=\sum_{i}b_{i}\left(d_{i}-b_{i}\right)$ (3)
    $f_{2}$ $=\sum_{i}\left(d_{\max}-d_{i}\right)\left(d_{i}-b_{i}\right)$ (4)
    $f_{3}$ $=\sum_{i}b_{i}^{2}\left(d_{i}-b_{i}\right)^{4}$ (5)
    $f_{4}$ $=\sum_{i}\left(d_{\max}-d_{i}\right)^{2}\left(d_{i}-b_{i}\right)^{4},$ (6)
    where $d_{\max}=\max d_{i}$. Points in $D$ with infinite death value are ignored.
    13) Complex Polynomial
    This vectorization computes the coefficients of a complex polynomial
    $C_{X}(z)=\prod_{i}\left[z-X\left(b_{i},d_{i}\right)\right],$ (7)
    where $X:R\to C$ is any of the following three functions:
    $R(x,y)$ $=x+iy,$ (8)
    $S(x,y)$ $=\begin{cases}\frac{y-x}{\sqrt{2(x^{2}+y^{2})}}(x+iy),&\textrm{if }(x,y)\neq(0,0)\\ 0&\textrm{otherwise}\end{cases}$ (9)
    Points in $D$ with infinite death value are ignored.
    15) Functional Data Analysis:
    $g(t)=\sum_{i}I_{[b_{i},d_{i})]}(t)=\sum_{k}c_{k}\xi_{k}(t)$